Choose two numbers X and Y independently at random from the unit interval [0,1] with the uniform density. The probability that

X^2+Y^2>0.81

is ?
Answer #1

for given data:

P( X^2+Y^2>0.81 ) =P(Area of the shaded region on the curve) =1-P(area of unshaded region )

=1-(/4)*r2 =1-(22/28)*0.812 =1-0.5153 =0.4847